### v3

# Correcting Nonresponse Bias Using Panel Data

Abstract
When subjects who respond to requests for data, such as in surveys or post-treatment follow-up, are not representative of the population as a whole, inferences drawn from the data can be misleading. We show that if subjects’ accumulated requests and responses over time are recorded and organized as panel data, requests can be used as instruments to correct for nonresponse bias even if total requests are not randomized between subjects. We demonstrate our method by estimating an 18-percentage-point gender gap in entrepreneurial career intentions using a survey of undergraduates at the University of Wisconsin-Madison with a 15% response rate and a 20-percentage-point intention gap among respondents.
JEL Codes: C83, L26, J16
Keywords: Survey Nonresponse, Attrition, Sample Selection, Entrepreneurship, Gender
1 Introduction
Researchers, governments, and private companies routinely request voluntary data submission, often more than once, from samples of populations of interest. These data gathering efforts usually only succeed with a portion of subjects. For example, this pattern occurs in experiments with delayed measurement of outcomes, such as those investigating effects of masks on COVID 19 transmission (abaluck2022impact), effects of education on STIs and fertility (duflo2015education), and effects of neighborhoods on health and safety (katz2001moving). The same pattern arises in survey data gathered by universities requesting student course reviews, by political pollsters seeking to predict election results, by private businesses seeking feedback on product satisfaction, and by governments seeking information from individuals, businesses, and nonprofit institutions. [^5] For instance, the Census Bureau (spiers2022strategic), the Bureau of Labor Statistics (nlsy97retention), and the Internal Revenue Service (irscp59) all document procedures for making multiple attempts to elicit responses from surveyed individuals for data they collect. If the individuals who voluntarily provide data are not representative of the population as a whole, inferences drawn from the data about the real world can be misleading.
An insight from heckman1976common; heckman1979sample is that variation in data observation rates between otherwise similar subpopulations can be used to correct for selection bias. To perform such a correction, it is necessary to retain data on nonrespondents along with respondents. Our core insight is that response behavior is observed for each subject each time they are sent a data request, rather than once. Retaining all such observations, at odds with standard practice, naturally provides variation in data observation rates between otherwise similar subpopulations. Briefly, we propose that researchers retain panel data on subjects’ responses at the time of each data request so that cumulative request counts can be used as instrumental variables to correct for nonresponse bias. [^6] Methods that leverage instruments include parametric methods such as those of heckman1979sample and the extension to binary outcomes of van1981demand, as well as nonparametric bounding methods such as those of manski1990nonparametric, manski1994selection, and manski2000monotone.
We join behaghel2015please and dutz2021selection in ordering subjects’ responsiveness by their response timings to correct for nonresponse bias. Our primary distinction from their methods is to use response and request timing to construct a panel dataset containing subjects’ responses as of each data request, rather than using response timing to order responsiveness between subjects in cross-sectional data. Constructing a panel dataset has two advantages. First, it makes explicit the assumption that potential responses are unaffected by time, which is made implicitly when potential responses are modeled cross-sectionally. Second, it converts response timing data into a format that accommodates existing nonresponse corrections that leverage instruments, as well as accommodating additional instruments that may or may not vary over time, such as randomized response incentives, without requiring revisions to estimation procedures.
Our main result is a proof that local average responses (LARs) for individuals who comply with each data request are identified from the joint distribution of responses and response rates for all requests. Identification of multiple local average responses is sufficient to estimate or bound population averages of requested variables using existing estimation methods that correct for selection bias, which are united by the marginal treatment response framework of heckman2005structural; heckman2007econometric. We establish identification under assumptions that resemble those of imbens1994identification and angrist1996identification. Identification follows regardless of whether requests are randomized or sent uniformly to all subjects if potential responses are time-invariant, and identification fails regardless of randomization if potential responses vary systematically with time.
In Section 3, we demonstrate the use of nonrandom request instruments by estimating the gender gap in entrepreneurial career intentions among undergraduates using survey data from the University of Wisconsin-Madison. Estimating a parametric van1981demand model for binary outcomes, we find an 18-percentage-point nonresponse-corrected male-female gap in entrepreneurial intentions. We find statistically insignificant evidence of positive selection for men and negative selection for women driving the 20 percentage point uncorrected gap among respondents. This model is identified with two requests, so we also implement and pass a “predictable trends” overidentification test of the null that later requests have no direct effect on the outcome, supporting the validity of our parametric and nonparametric assumptions in this application. In Appendix B, we estimate labor market outcomes using a synthetic version of the Norway in Corona Times survey. We construct our synthetic dataset to match respondent mean estimates reported by dutz2021selection, which are sufficient to estimate population means with our method. The 95% confidence intervals of our estimated population means cover ground truth averages from administrative data reported by these authors for five out of six survey variables, with respondent averages differing significantly from the ground truth for all variables.
2 Correcting Nonresponse With Multiple Requests
We consider a panel of $N$ subjects indexed by $i$ who are observed in time periods t=0,1,,Tt=0,1,...,T, with data collection commencing when $t=1$ and concluding when $t=T$. Subject $i$ receives data requests for an unobserved time-invariant requested variable, denoted by $Y_{i}^{*}$, with their observed requests accumulated by the end of time $t$ given by $R_{it}$. In each time period, we also observe their response choice $S_{it}$, with $S_{it}=1$ if they respond during time $t$ and $S_{it}=0$ if they do not, with their response, $Y_{it}$, observed if and only if $S_{it}=1$. We assume $S_{it}=1$ for subject $i$ at most once. Finally, we denote subject $i$’s retained response choice for time $t$ as $\hat{S}_{it}=\sum_{k=0}^{t}S_{ik}$, with their retained response for time $t$ given by $\hat{Y}_{it}=\sum_{k=0}^{t}Y_{ik}S_{ik}$.
2.1 Bias in Selected Samples
In settings with nonresponse, the sample mean of retained responses in period $T$ among respondents is a consistent estimate of $\mathbb{E}[\hat{Y}_{iT}|\hat{S}_{iT}=1]$, with expected bias relative to $\mathbb{E}[Y_{i}^{*}]$ given by
$B_{w}=\mathbb{E}[\hat{Y}_{iT}|\hat{S}_{iT}=1]-\mathbb{E}[Y_{i}^{*}].$
With random nonresponse and no measurement error, $\mathbb{E}[\hat{Y}_{iT}|\hat{S}_{iT}=1]=\mathbb{E}[Y_{i}^{*}]$, and valid inferences on populations are possible using sample means of observed responses. If these conditions are not met, identification of population averages requires additional information.
2.2 Modeling Nonresponse
We assume that observed variables are realizations of potential outcomes, following rubin1974estimating. First, let $S_{it}=S_{it}(R_{it})\prod_{j=0}^{t-1}(1-S_{ij}(R_{ij}))$ where $S_{it}(r)\in\{0,1\}$ indicates subject $i$’s willingness to respond to $r$ requests at time tt. [^7] We use this definition in lieu of $S_{it}=S_{it}(R_{it})$ because we assume that subjects respond at most once. Similarly, let $Y_{it}=Y_{it}(S_{it},R_{it})=Y_{it}(1,R_{it})S_{it}+Y_{it}(0,R_{it})(1-S_{it})$ with $Y_{it}(s,r)$ denoting the response subject $i$ would give at time $t$ if they made response choice $s$ given $r$ accumulated requests, with $Y_{it}(0,r)$ missing for all $i$, $t$, and $r$. $S_{it}(r)$ and $Y_{it}(s,r)$ are defined for all values in the supports of their domains, $r\in\{0,...,\max(R_{it})\}$ and $s\in\{0,1\}$, respectively, while they are observed only for their realized values.
Following the marginal treatment response framework of heckman2005structural; heckman2007econometric, we assume that selection bias is driven by dependence between responses and response aversion. Specifically, we define the marginal survey response function $m(u)\equiv\mathbb{E}[Y_{i}^{*}|U_{it}=u]$ where $U_{it}\in[0,1]$ denotes the response aversion percentile of subject $i$ at time $t$. We assume that response aversion determines potential response choices such that
$S_{it}(r)=1(U_{it}\leq P(r)),$ (1)
for all $r$, where $P(r)=\mathbb{E}[S_{it}(r)]$ gives the willing response propensity for request $r$. We then have
$\mathbb{E}[Y_{i}^{*}|P(r^{\prime})<U_{it}\leq P(r)]=\mathbb{E}[Y_{i}^{*}|S_{it}(r)-S_{it}(r^{\prime})=1],$ (2)
for all $r$ and $r^{\prime}$.
It follows that we can learn about $m(u)$ if we can estimate $\mathbb{E}[Y_{i}^{*}|S_{it}(r)-S_{it}(r^{\prime})=1]$, which we refer to as the local average response for $r-r^{\prime}$ compliers. If $m(u)$ is parameterized, as in a heckman1979sample model, point estimation is possible if there are weakly more identified values of $P(r)$ than parameters. Otherwise, $m(u)$ can be bounded using methods such as those of horowitz2000nonparametric or lee2009training. Local average responses are identified as
$\mathbb{E}[Y_{i}^{*}|S_{it}(r)-S_{it}(r^{\prime})=1]=\frac{\mathbb{E}[\hat{Y}_{it}|R_{it}=r]-\mathbb{E}[\hat{Y}_{it}|R_{it}=r^{\prime}]}{\mathbb{E}[\hat{S}_{it}|R_{it}=r]-\mathbb{E}[\hat{S}_{it}|R_{it}=r^{\prime}]}$ (3)
if requests are valid instruments for requests, so we proceed by establishing this result with assumptions similar to those used to establish instrument validity by imbens1994identification and angrist1996identification, which were shown by vytlacil2002independence to imply the relationship in (1).
2.3 Local Average Response Identification Assumptions
Our goal is to attribute differences in $\hat{Y}_{it}$ across different values of $R_{it}$ to the average of $Y_{i}^{*}$ for compliers who require more requests to induce response. If more requests are systematically sent to subjects or time periods with unrepresentative potential responses, we cannot reliably attribute differences in responses between requests to marginal compliers. We begin by assuming that requests are sent in such a way that subjects’ potential responses and potential response choices across time are unrelated to their request histories.
Condition 1 (Intertemporal Independence):
[…]
We next assume that requests do not directly affect potential responses, allowing us to write $Y_{it}(1)=Y_{it}(1,r)$.
Condition 2 (Exclusion):
$Y_{it}(1,r)=Y_{it}(1,r^{\prime})$ for all $r,r^{\prime}$, for all $i$ and $t$.
The importance of avoiding experimenter demand effects or priming in surveys is well established (stantcheva2023run). Our assumption expresses this as a special case of similar assumptions made in the context of sample selection by heckman1979sample and in the context of treatment effect identification by angrist1996identification. In contrast with intertemporal independence, exclusion is more likely to be violated due to continuous time variation in potential responses and requests if time periods are defined coarsely, as subjects receiving different numbers of requests in the same discrete time period may receive them at different points in continuous time. [^9] Following imbens1994identification, we’ve assumed implicitly that $Y_{it}(s,r)$ and $S_{it}(r)$ are fixed for all $s,r$, for all $i,t$ (SUTVA). This assumption is consequentially violated if potential responses or potential response choices systematically vary in continuous time within discrete time periods.
With independence and exclusion, differences in responses between requests can only be explained by differences in responses between distinct complier groups, defined by common values of $S_{it}(r)$ for all $r$. Our next assumption restricts the possible number of these groups by imposing a natural ordering on them.
Condition 3 (Intertemporal Monotonicity):
a. $S_{it}(r)\geq S_{it}(r^{\prime})$ for all $r>r^{\prime}$, for all $i$ and $t$.
b. $S_{it}(r)=S_{it^{\prime}}(r)$ for all $t$ and $t^{\prime}$, for all $i$ and $r$.
c. $R_{it}\geq R_{it^{\prime}}$ for all $t>t^{\prime}$, for all $i$.
d. $R_{i0}=0$ for all $i$ and $S_{it}(0)=0$ for all $i$ and $t$.
Our assumption strengthens that of imbens1994identification by imposing a sign restriction on the effects of requests on response choices, time-invariance on subjects’ potential response choices, and intuitive restrictions on the accrual of requests and responses over time. Our monotonicity assumption implies that $S_{i}(R_{it})=\hat{S}_{it}$, which enables identification of $P(r)$ and $\mathbb{E}[Y_{it}(1,R_{it})|S_{i}(R_{it})=1]$ from average retained response choices and retained responses. Relatedly, we assume that requests strictly increase response rates, slightly strengthening the first part of our monotonicity assumption.
Condition 4 (Relevance):
$P(r)>P(r^{\prime})$ for all $r>r^{\prime}$.
As in treatment effects applications, relevance can be tested by estimating $P(r)$ for all $r$, with $P(r)=\mathbb{E}[\hat{S}_{it}|R_{it}=r]$ if intertemporal independence and intertemporal monotonicity hold.
Independence, exclusion, and monotonicity allow us to determine local average responses for compliers to each request. We also assume that responses are free of systematic measurement error.
Condition 5 (Classical Measurement Error):
$\mathbb{E}[Y_{it}(1,r)S_{it}(r)]=\mathbb{E}[Y_{i}^{*}S_{it}(r)]$ for all $r$.
This assumption is weaker than $Y_{it}(1,r)=Y_{i}^{*}$, which is common in the related literature, because it allows for idiosyncratic shocks that cancel out within complier groups, such as if students’ potential responses regarding career intentions change over time in response to new information, such as grades on assignments. [^10] We hope that our consideration of measurement error will forestall inaccurate presumptions that weaker assumptions, such as $\mathbb{E}[Y_{it}(1,r)]=\mathbb{E}[Y_{i}^{*}]$, are sufficient. Intertemporal independence and exclusion rule out systematic time-varying measurement error and measurement error caused by requests, respectively, but they do not rule out systematic inaccuracies within or across complier groups. For example, if less conscientious subjects give systematically inaccurate responses to late requests, nonresponse corrections using request instruments will tend to replicate similar inaccuracy in their predictions for non-respondents.
2.4 Local Average Response Identification Proof
Our main result proves that local average responses are identified under conditions 1–5.
Theorem 1:
If Conditions 1–5 hold, the local average response for subjects who respond to request $r$ but not $r^{\prime}$ is identified as
$\mathbb{E}[Y_{i}^{*}|S_{i}(r)-S_{i}(r^{\prime})=1]=\frac{\mathbb{E}[\hat{Y}_{it}|R_{it}=r]-\mathbb{E}[\hat{Y}_{it}|R_{it}=r^{\prime}]}{\mathbb{E}[\hat{S}_{it}|R_{it}=r]-\mathbb{E}[\hat{S}_{it}|R_{it}=r^{\prime}]}$
from the joint distribution of $Y$, $S$, and $R$ for any $r>r^{\prime}$ such that $\mathbb{E}[\hat{Y}_{it}|R_{it}=r]$ and $\mathbb{E}[\hat{Y}_{it}|R_{it}=r^{\prime}]$ are finite and $P(r)>P(r^{\prime})$.
Proof.
Assuming that Conditions 1–3 and 5 hold, we have
$\mathbb{E}[\hat{Y}_{it}|R_{it}=r]$ $=\mathbb{E}[\sum_{k=0}^{t}Y_{ik}(1,R_{ik})S_{ik}(R_{ik})\prod_{j=0}^{k-1}(1-S_{ij}(R_{ij}))|R_{it}=r]$
$=\mathbb{E}[\sum_{k=1}^{t}Y_{ik}(1,R_{ik})(S_{i}(R_{ik})-S_{i}(R_{i,k-1}))|R_{it}=r]$
$=\mathbb{E}[Y_{it}(1)\sum_{k=1}^{t}(S_{i}(R_{ik})-S_{i}(R_{i,k-1}))|R_{it}=r]$
$=\mathbb{E}[Y_{it}(1)S_{i}(r)]$
$=\mathbb{E}[Y_{i}^{*}S_{i}(r)].$
The first equality rewrites retained responses in terms of potential outcomes. The second equality follows from monotonicity. The third equality follows from independence and exclusion. The fourth equality follows from monotonicity and independence. The fifth equality follows from classical measurement error.
Assuming that conditions 1–5 hold, it follows that the instrumental variables estimand identifies the local average response for $r>r^{\prime}$ as
$\frac{\mathbb{E}[\hat{Y}_{it}|R_{it}=r]-\mathbb{E}[\hat{Y}_{it}|R_{it}=r^{\prime}]}{\mathbb{E}[\hat{S}_{it}|R_{it}=r]-\mathbb{E}[\hat{S}_{it}|R_{it}=r^{\prime}]}$ $=\frac{\mathbb{E}[Y_{i}^{*}S_{i}(r)]-\mathbb{E}[Y_{i}^{*}S_{i}(r^{\prime})]}{P(r)-P(r^{\prime})}$
$=\mathbb{E}[Y_{i}^{*}|S_{i}(r)-S_{i}(r^{\prime})=1],$
where the first equality makes the substitution $\mathbb{E}[\hat{Y}_{it}|R_{it}=r]=\mathbb{E}[Y_{i}^{*}S_{i}(r)|R_{it}=r]$ as derived above, and makes the substitution $\mathbb{E}[\hat{S}_{it}|R_{it}=r]=P(r)$ following the same steps. The second equality follows from monotonicity and relevance. ∎
Our result establishes conditions under which data requests are valid instruments for retained responses in data where subjects’ requests and responses are observed multiple times. It follows that established methods that use instrumental variables to address selection bias, such as that of heckman1979sample, can be used in settings without explicitly randomized response incentives. The insight underlying these methods is that differences in responses between otherwise-similar subjects with different response rates can be explained by dependence between potential responses and response aversion. A more thorough review of these methods is provided by dutz2021selection, along with a generalized model of survey aversion with heterogeneous effects of instruments on responsiveness.
3 Application: Gender Entrepreneurship Gap
We use our method to estimate the gender gap in entrepreneurial aspirations among undergraduate students. There is a large gender gap in entrepreneurship among working-age adults (aldrich2005entrepreneurship), perhaps driven at least in part by a gender gap in early-stage funding for new ventures (canning2012women; greene2003women). Our paper contributes to this literature by investigating pre-market gender gaps in entrepreneurial career interest, which may be more likely to develop due to intrinsic interest gaps or pre-market discrimination rather than labor market discrimination. This application showcases the importance of researcher judgments in defining the time index, adjusting for nonrandom variation in requests, and testing modeling assumptions in cases where there are more requests than are needed for identification. In Appendix B, we demonstrate our method’s accuracy by showing that nonresponse-corrected estimates are closer to ground truth values from administrative data than respondent averages for the Norway in Corona Times Survey, using survey and ground truth averages reported by dutz2021selection.
3.1 Data
We use a survey of the entrepreneurial intentions of the undergraduate population at University of Wisconsin–Madison that was implemented every fall from 2015 to 2022, as well as in spring 2020 and 2021. Students could choose between “Yes,” “No,” or “I Don’t Know,” when responding to a question regarding whether they intend to pursue a career in entrepreneurship. We code an “Intention” variable equal to 1 if the student answered “Yes” and 0 if they answered “No” or “I Don’t Know.” We observe term-specific survey request stratum indicators for groups of students who are sent requests at the same time through survey software. We observe timestamps, in seconds, for each request made to each stratum and for each student’s response if they respond. Students who opted out of the survey in a prior year have no stratum indicators and no request or response timestamps in the raw data, whereas students who respond to early requests have stratum indicators and request timestamps for all intended requests in raw data, though they did not receive additional actual requests after responding.
Overall, 103,536 unique students and 333,201 total student-terms were surveyed. We merge the survey data to administrative records containing information on students regardless of whether they responded to the survey. Business Major and STEM Major are set to one if the student has any major in the term on the business school’s list of majors or on the list of STEM majors provided by ice2020, respectively. GPA refers to the student’s cumulative grade point average that varies across terms, ACT Math is the average of students’ ACT math test scores on file, and ACT Verbal is the average of the sum of English and Reading portions of the ACT test. [^11] We convert SAT scores into ACT scores for students without ACT scores using the 2018 ACT/SAT concordance tables provided by act.org. Female, Racial Minority, and International are binary indicators for female gender from binary gender information in raw data, listing a race other than white, and having a recorded country of origin other than the U.S., respectively. We drop 6,075 student-terms with missing GPAs (students who dropped out in their first term) and 33,296 observations with missing ACT scores, leaving us with 290,588 student-terms and 82,000 unique students. There were 42,902 total survey responses, with a response rate of approximately 14% for men and 16% for women. Sample means of key variables broken down by response timing are provided in Table 1.
Table 1: Summary Statistics
Early Respondents Late Respondents Nonrespondents
Men Women All Men Women All Men Women All
(1) (2) (3) (4) (5) (6) (7) (8) (9)
Intention 0.378 0.169 0.261 0.366 0.180 0.266
Female 0.000 1.000 0.561 0.000 1.000 0.538 0.000 1.000 0.513
Business Major 0.126 0.090 0.106 0.127 0.096 0.110 0.120 0.082 0.100
STEM Major 0.492 0.305 0.387 0.488 0.297 0.385 0.464 0.288 0.374
GPA 3.427 3.450 3.440 3.413 3.450 3.433 3.287 3.398 3.344
ACT Math 29.813 27.638 28.594 29.827 27.515 28.584 29.393 27.153 28.245
ACT Verbal 56.956 56.929 56.941 57.349 57.021 57.173 56.287 56.125 56.204
Racial Minority 0.244 0.228 0.235 0.287 0.262 0.274 0.304 0.283 0.293
International 0.100 0.065 0.080 0.110 0.071 0.089 0.106 0.072 0.089
Observations 10154 12958 23112 9147 10643 19790 120730 126956 247686
Notes: Variable means for men and women by responder group. Early Respondents respond prior to the second data request, while Late Respondents respond any time after the second request.
In the raw data, men’s entrepreneurial intention rate drops from 37.8% among early respondents to 36.6% among late respondents, while women’s entrepreneurial intention rate rises from 16.9% among early respondents to 18.0% among late respondents. Most of the other variables we review do not exhibit any strong pattern, with the exception of racial minority status. We find that racial minorities are underrepresented among survey respondents, representing 23.5% of early respondents, 27.4% of late respondents, and 29.3% of nonrespondents.
3.2 Constructing a Request-Response Panel
Our results from Section 2 require panel data wherein subjects are observed multiple times, so we construct such a panel. In our setting, students who opted out of the survey in a prior term and students who respond early in a given term do not receive subsequent requests. It follows that actual requests almost surely violate the independence assumption in Section 2, as students receiving each number of requests differ systematically. We address this by randomly assigning opt-outs to strata ex post and assuming $S_{i}(r)=0$ for all $r$ for these students, then imputing intended request timestamps for each student using the request timestamps of nonrespondents in their stratum.
For each student-term in a stratum, we produce observations for $t$ in $0,1,...,R_{i}$ such that time periods are defined as intervals in which a subject has received a given number of intended requests, with $R_{it}=t$ for all $i$ and $t$. Independence and monotonicity require that the times at which requests were made had representative potential responses and that subjects had sufficient time to respond between requests, respectively. We report request timestamps in Appendix A, revealing that the shortest interval between requests is three days, which we assume was adequate time for students to respond. We set $S_{it}=1$ if subject $i$’s response timestamp is between those defining the beginning and end of period $t$, with $Y_{it}$ given by their response to the survey if $S_{it}=1$. We forward-impute retained response choices $\hat{S}_{it}$ and retained responses $\hat{Y}_{it}$ as described in Section 2.
3.3 Nonresponse-Corrected Entrepreneurial Intentions by Gender
(a) Panel A: Men
(b) Panel B: Women
Figure 1: Observed and Estimated Entrepreneurial Intention, by Gender
Notes: Average entrepreneurial intention rates as a function of response aversion. Local average responses are given by $\beta$ from the equation $\hat{Y}_{it}=\hat{S}_{it}\beta+\epsilon_{i}$, which we estimate via 2SLS using $I(R_{it}=r)$ as an instrument for $\hat{S}_{it}$ in the subsample with $R_{it}\in[r-1,r]$, for $r=1,2$. Corrected estimates by survey aversion percentile are obtained by estimating a van1981demand model and calculating $\hat{m}(u)=\Phi((\hat{\beta}+\hat{\rho}\Phi^{-1}(1-u))/\sqrt{1-\hat{\rho}^{2}})$ for $u\in[0,1]$. Standard errors in parentheses are clustered at the student level.
We begin by estimating the raw gender gap in entrepreneurial intentions without adjusting for the covariates listed in Table 1. Because requests vary in number from two to seven between terms, we drop all observations with $R_{it}>2$ to support our independence assumption for the current specification. We estimate a parametric van1981demand selection model separately by gender via maximum likelihood, in which
$\hat{S}_{it}=\begin{cases}&1\mbox{ if }Z_{it}\alpha+u_{i}\geq 0,\\ &0\mbox{ otherwise,}\end{cases}$ (4)
$\hat{Y}_{it}=\begin{cases}1\mbox{ if }X_{i}\beta+\epsilon_{i}\geq 0\quad\&\quad&\hat{S}_{it}=1,\\ 0\mbox{ if }X_{i}\beta+\epsilon_{i}<0\quad\&\quad&\hat{S}_{it}=1,\\ \cdot\mbox{ if }&\hat{S}_{it}=0,\end{cases}$ (5)
and
[…]
where $X_{i}=1$, $Z_{it}=[X_{i},I(R_{it}=2)]$, and $\rho$ gives the correlation between outcome residuals $\epsilon_{i}$ and response preference $u_{i}=\Phi^{-1}(1-U_{i})$, where $U_{i}$ is $i$’s response aversion percentile in the notation of Section 2 and $\Phi^{-1}(\cdot)$ denotes the inverse of the normal CDF. [^12] Our 15% response rate is insufficient for informative bounds from conservative approaches such as that of horowitz2000nonparametric, and we doubt that gender satisfies the monotonicity assumption of lee2009training or behaghel2015please, rendering the “local gender gap” among survey respondents unidentified, to say nothing of its policy-relevance.
This model implies that expected intention as a function of response aversion is $m(u)=\Phi((X_{i}\beta+\rho\Phi^{-1}(1-u))/\sqrt{1-\rho^{2}})$, with the expectation conditional only on $X_{i}$ given by $\int_{0}^{1}m(u)du=\Phi(X_{i}\beta)$. We graphically present our estimates of $m(u)$ for men and women in Figure 1 for $u\in[0,1]$. Local average responses and sample averages for compliers to the first two requests are shown for reference. The levels of entrepreneurial intention for early and late respondents show a decreasing pattern for men and an increasing pattern for women, as in Table 1. Differences in selection bias between men and women, reflected by their $\rho$ estimates, explain the difference between the 20 percentage point uncorrected gap estimate and the 12 percentage point corrected gap estimate.
3.4 Oaxaca-Blinder-Kitagawa Decomposition
In addition to estimating average entrepreneurial intention for men and women, we also decompose the gender intention gap using a nonlinear extension of methods developed by kitagawa1955components, oaxaca1973male, and blinder1973wage. Our decomposition estimates, for each covariate in $X_{i}$, the extent to which the gender gap would be reduced if women’s average value of that single variable were equal to that of men, and how much it would be reduced if that variable’s coefficient for women were equal to the coefficient for men. These results are shown in Table 2.
To implement the decomposition, we estimate the model in (4)-(6) jointly for men and women via maximum likelihood using gender-specific parameters $(\beta_{m},\alpha_{m},\rho_{m})$ for men and $(\beta_{w},\alpha_{w},\rho_{w})$ for women. To avoid overweighting terms in which more requests were sent, which have more observations, we weight each observation by the inverse of the total number of requests sent to that student in that term. We define $X_{i}$ as all covariates listed in Table 1 along with term fixed effects, and we define $Z_{it}$ as binary indicators for each value in $R_{it}=1,2...,7$ interacted with $X_{i}$.131313blandhol2022tsls recommend rich covariate interactions for instrumental variables when estimating local average treatment effects via two stage least squares. The estimator and target parameters are different here, but we find their argument convincing that if model assumptions hold conditional on $X_{i}$, then $X_{i}$ should be flexibly interacted with instruments to support identification. We order the sample of $N_{w}$ women and $N_{m}$ men such that subjects with $i=1,...,N_{w}$ are women and those with $i=N_{w}+1,...,N$ are men, with $N=N_{w}+N_{m}$.
Table 2: Oaxaca-Blinder-Kitagawa Decomposition of the Entrepreneurial Intention Gender Gap
Uncorrected Corrected
Men Women Gap Men Women Gap
Variables (1) (2) (3) (4) (5) (6)
Intention 0.381 0.179 0.202 0.375 0.192 0.183
(0.005) (0.003) (0.005) (0.020) (0.015) (0.025)
X
Business Major 0.120 0.084 0.004 0.120 0.084 0.004
(0.001) (0.001) (0.000) (0.001) (0.001) (0.000)
STEM Major 0.468 0.290 -0.005 0.468 0.290 -0.006
(0.001) (0.001) (0.001) (0.001) (0.001) (0.001)
GPA 3.305 3.406 0.005 3.305 3.406 0.005
(0.001) (0.001) (0.001) (0.001) (0.001) (0.001)
ACT Math 29.452 27.221 0.003 29.452 27.221 0.003
(0.011) (0.010) (0.002) (0.011) (0.010) (0.002)
ACT Verbal 56.405 56.258 -0.000 56.405 56.258 -0.001
(0.023) (0.022) (0.000) (0.023) (0.022) (0.000)
Racial Minority 0.299 0.277 0.001 0.299 0.277 0.001
(0.001) (0.001) (0.000) (0.001) (0.001) (0.000)
International 0.105 0.072 0.005 0.105 0.072 0.005
(0.001) (0.001) (0.000) (0.001) (0.001) (0.001)
All X 0.013 0.013
(0.002) (0.003)
$\beta$
Business Major 0.449 0.457 -0.000 0.439 0.457 -0.000
(0.036) (0.038) (0.001) (0.037) (0.039) (0.002)
STEM Major -0.039 -0.118 0.005 -0.041 -0.122 0.006
(0.024) (0.027) (0.003) (0.025) (0.028) (0.003)
GPA -0.181 -0.198 0.015 -0.186 -0.198 0.010
(0.024) (0.026) (0.031) (0.026) (0.027) (0.034)
ACT Math 0.007 0.006 0.012 0.008 0.006 0.014
(0.004) (0.004) (0.039) (0.004) (0.004) (0.042)
ACT Verbal -0.013 -0.013 0.008 -0.013 -0.013 0.010
(0.002) (0.002) (0.034) (0.002) (0.002) (0.037)
Racial Minority 0.122 0.191 -0.006 0.105 0.182 -0.007
(0.032) (0.030) (0.004) (0.034) (0.031) (0.004)
International 0.379 0.590 -0.006 0.403 0.609 -0.006
(0.049) (0.053) (0.002) (0.051) (0.055) (0.002)
All $\beta$ 0.030 0.030
(0.042) (0.047)
All Unexplained Gaps 0.159 0.140
(0.043) (0.059)
Auxiliary Parameters
$\rho$ 0 0 0 0.010 -0.032 0
(.) (.) (.) (0.034) (0.035) (.)
Log-Likelihood -67,349 -74,546 -141,895 -57,558 -64,508 -122,066
Observations 140,031 150,557 290,588 604,031 652,290 1,256,321
Model Tests Men Women Both Gap
p-value: $\mathbb{E}[\mathbb{E}[Y_{i}^{*}|S_{i}=1,X_{i}]]=\mathbb{E}[\mathbb{E}[Y_{i}^{*}|X_{i}]]$ 0.743 0.399 0.664 0.449
p-value: $\rho=0$ 0.762 0.363 0.631 0.387
p-value: $\beta_{R}=0$ (no request effects) 0.135 0.411 0.200
Notes: Uncorrected estimates obtained by estimating a van1981demand model using only the final request observation for each student-term with the constraint $\rho=0$. Corrected estimates obtained by estimating the same model using all student-term-request observations with unconstrained $\rho$, with observations weighted by the inverse of total requests by student-term. Group-specific population mean estimates of variables and parameters, as indicated, are in columns (1), (2), (4), and (5), with sample means of $X$ calculated across student-term observations. Columns (3) and (6) give uncorrected and corrected estimated contributions of each variable or coefficient to the estimated gap in entrepreneurial intentions, using equations (7), (8), (9), (10), and (11). $\mathbb{E}[\mathbb{E}[Y_{i}^{*}|S_{i}=1,X_{i}]]$ is estimated as the average of $\Phi(X_{i}\hat{\beta}^{uncorrected})$ and $\mathbb{E}[\mathbb{E}[Y_{i}^{*}|X_{i}]]$ is estimated as the average of $\Phi(X_{i}\hat{\beta}^{corrected})$ over student-term observations. Term fixed effects, included in $X_{i}$ to support the independence assumption, are treated as part of the unexplained gap. Standard errors clustered at the individual level are in parentheses.
We estimate the contribution of $x_{i}\subset X_{i}$ to the entrepreneurial intention gap as
$\hat{\Delta}_{x}=\sum_{i=1}^{N_{w}}\frac{\Phi(X_{i}\hat{\beta}_{w}+(\bar{x}_{m}-\bar{x}_{w})\hat{\beta}_{w,x})-\Phi(X_{i}\hat{\beta}_{w})}{N_{w}},$ (7)
wherein we denote the sample average of $x_{i}$ for men and women as $\bar{x}_{m}$ and $\bar{x}_{w}$, respectively, and we denote male and female coefficients on $x_{i}$ as $\beta_{m,x}$ and $\beta_{w,x}$, respectively. [^14] Fairlie’s (fairlie2005extension) decomposition for nonlinear models considers the case in which the marginal distribution of $x_{i}$ for women is set equal to men’s, while ours considers the case in which the mean of $x_{i}$ for women is set equal to men’s via a level shift of women’s distribution. We view the counterfactual we consider as likely simpler to understand and simpler to implement for policymakers, though describing $\beta_{w,x}$ as the “effect” of changing $x_{i}$ for women requires assumptions beyond those that either we or fairlie2005extension make. We similarly estimate the contribution to the average gender gap of $x_{i}$’s differential effect on $Y_{i}$ between men and women as
$\hat{\Delta}_{\beta_{x}}=\sum_{i=1}^{N_{w}}\frac{\Phi(X_{i}\hat{\beta}_{w}+x_{i}(\hat{\beta}_{m,x}-\hat{\beta}_{w,x}))-\Phi(X_{i}\hat{\beta}_{w})}{N_{w}}.$ (8)
In addition to estimating the contribution of each variable and its coefficient on the gender-intention gap, we also estimate the contribution of the entire vector $X_{i}$ on the gap as
$\hat{\Delta}_{X}=\sum_{i=1}^{N_{w}}\frac{\Phi(X_{i}\hat{\beta}_{w}+(\bar{X}_{m}-\bar{x}_{w})\hat{\beta}_{w})-\Phi(X_{i}\hat{\beta}_{w})}{N_{w}},$ (9)
defining the vector of the means of all elements in $X$ for people with $W_{i}=w$ as $\bar{X}_{w}$. Similarly, we estimate the contribution of the entire vector of gender differences in coefficients on the gender-intention gap as
$\hat{\Delta}_{\beta}=\sum_{i=1}^{N_{w}}\frac{\Phi(X_{i}\hat{\beta}_{w}+X_{i}(\hat{\beta}_{m}-\hat{\beta}_{w}))-\Phi(X_{i}\hat{\beta}_{w})}{N_{w}}.$ (10)
We estimate the remaining gap, which is driven by the interaction of male-female differences in $X$ and male-female differences in $\beta$, distributional differences in $X$ between men and women, and nonlinearity of the normal distribution as
$\hat{\Delta}_{R}=\sum_{i=N_{w}+1}^{N}\frac{\Phi(X_{i}\hat{\beta}_{m})}{N_{m}}-\sum_{i=1}^{N_{w}}\frac{\Phi(X_{i}\hat{\beta}_{w})}{N_{w}}-\hat{\Delta}_{X}-\hat{\Delta}_{\beta}.$ (11)
Table 2 contains estimated means of the variable or coefficient listed in each row for men, women, and the contribution of that item to the intention gap, using predicted values from a naive probit in columns 1–3, and using our proposed selection correction model in columns 4–6. $X$ is observed for all individuals regardless of survey response, so the means of $X$ by gender are the same for the uncorrected model and the corrected model. Columns 3 and 6 report estimated contributions of each row quantity to the gender gap as described in equations (7), (8), (9), (10), and (11). We perform our decomposition for all variables in Table 1, with term fixed effects included in $X_{i}$ contributing only to the unexplained gap.
We estimate the overall gender gap by calculating
$\hat{\Delta}=\sum_{i=N_{w}+1}^{N}\frac{\Phi(X_{i}\hat{\beta}_{m})}{N_{m}}-\sum_{i=1}^{N_{w}}\frac{\Phi(X_{i}\hat{\beta}_{w})}{N_{w}},$ (12)
where our estimates of $(\hat{\beta}_{w},\hat{\beta}_{m})$ differ by estimation method. The estimated gender-intention gap from our preferred specification is 18.3 percentage points, with a naive probit producing an estimate of 20.2 percentage points. Differences between the corrected and uncorrected model are due to the inclusion of the $\rho$ parameter in the corrected model (with $\rho=0$ fixed in the uncorrected model) which is positive if early respondents have higher intention and negative if they have lower intention. Our estimates of $\rho$ in the current specification are insignificant, suggesting limited nonresponse bias, though we see the same pattern with the signs of our point estimates here as we did in the no-controls specification shown in Figure 1. Along these same lines, our richer specification repeats the finding in Figure 1 that the uncorrected gap point estimate is larger than the corrected estimate, but here we fail to reject that they are the same in the population (p=0.449).
The difference in corrected estimates in Section 3.3 and the corrected estimates in Table 2 is due to the inclusion of controls and observations for requests three through seven, which were observed in some years but not others. The richer nonresponse correction specification in this section approximates a nonparametric specification that would fit a curve through local responses as in Figure 1 for each value of $X_{i}$, where the slope of the curve is determined by a gender-specific $\rho$ that is held fixed across $X_{i}$. For both men and women, $\rho$ is attenuated in the richer model, suggesting that observed covariates account for both late responses and differences in intention between late and early responses. Nonetheless, the same general pattern from Figure 1 persists, with positive selection for men ($\hat{\rho}_{m}>0)$ and negative selection for women $(\hat{\rho}_{f}<0)$, though these estimates are not statistically significant in this specification.
Our estimates are only reliable if the assumptions in Section 2 and equation (6) hold for this application. Under these assumptions, all local average responses should be explained by the marginal survey response function. We assess this by performing a “predictable trends” overidentification test wherein we include indicators for requests $R_{it}=3,...,7$ in $X_{i}$ with coefficient vector $\beta_{R}$, relying on requests one and two for identification. We fail to reject the null that later requests have no effect on responses at conventional levels for men (p=0.135), for women (p=0.411), and for both jointly (p=0.200). However, these point estimates are narrowly insignificant and mostly positive for men as shown in Figure 2, suggesting that later requests drive corrected point estimates upward toward the uncorrected estimates for men when they are included, increasing the estimated gender gap.
Figure 2: Event Study of Group Means Relative to Model Predictions, by Gender
Notes: We depict the probit coefficient point estimates and 95% confidence intervals for gender-specific estimated effects of requests $R_{it}>2$ on intention. We fail to reject that all effects are zero, with p=0.135 for men, p=0.411 for women, and p=0.200 for men and women jointly.
Most of the entrepreneurial intention gender gap is unexplained by our observed variables. While the contribution of differences in observed covariates to the gender gap of 1.3 percentage points is statistically significant at conventional levels, it explains a small fraction of the overall gap. Meanwhile, the total corrected contribution of gender differences in $\beta$ to the gender gap is larger at 3 percentage points, but it is not significant at conventional levels. The largest statistically significant individual contributors to the gender intention gap are gender differences in the rates of choosing a STEM major and gender differences in the relative rates of intention between STEM and non-STEM students. Interestingly, these factors approximately cancel each other out. On one hand, there are more male STEM majors, who have somewhat lower entrepreneurial intention than male non-STEM majors. Meanwhile, there are fewer female STEM majors, but female STEM majors have much lower entrepreneurial intention than female non-STEM majors.
Results from our preferred specification suggest a large gender gap in undergraduate entrepreneurial intention, with minimal nonresponse bias in this survey. Despite the apparent lack of significant nonresponse bias, our corrected estimates support substantially different inferences than the uncorrected estimates because of their larger standard errors. Perhaps the clearest evidence of this is that the uncorrected gap estimate strongly rejects a true population average equal to the corrected point estimate. Given that our parametric correction gives substantially more precise estimates than conservative approaches such as that of horowitz1998censoring, we recommend against drawing inferences while assuming no nonresponse bias in applications such as ours.
4 Conclusion
We have established that data requests made to subjects are valid nonresponse correction instruments in panel data containing their accumulated requests and responses over time, under assumptions similar to those routinely made when analyzing survey data. Specifically, we assume that each data request must be sent to a representative sample of the population, requests should not directly affect responses, and average responses to any given request should be free of systematic measurement error. While these assumptions are strong, their importance for securing identification is not specific to our method. In addition to these assumptions, we make a monotonicity assumption similar to that of imbens1994identification which allows us to order subjects by their response timing and extrapolate from respondents to nonrespondents using existing nonresponse correction methods.
Our procedure for constructing request instruments imposes minimal challenges for data collection and analysis. It requires no knowledge of econometrics by data collection administrators and no financial incentives for subjects, which can bias responses if subjects maximize their effective wages by minimizing time spent on the survey, such as by using artificial intelligence to concoct responses (westwood2025). Meanwhile, it naturally reveals rich variation in response rates when multiple reminders are sent, enabling model validation tests that compare late responses to their predicted values implied by early responses. These features enable routine use of our method when analyzing data obtained by contacting subjects, either via surveys or via similar procedures such as gathering lab data after medical experiments.
We illustrate our method by estimating the gender gap in entrepreneurial intentions among undergraduate students at the University of Wisconsin-Madison. To do so, we construct a panel dataset that contains a row for each student for each data request they receive, with their accumulated requests recorded in all rows and their responses recorded only in rows representing time periods after they respond. We find a large gender gap both without correcting for nonresponse and when correcting for nonresponse using requests as instruments in a van1981demand model, with the uncorrected estimates strongly rejecting the point estimate of the average gender gap obtained when correcting for nonresponse. In an additional application in Appendix B, we accurately and precisely estimate ground truth averages from administrative data for five out of six labor market indicators using only aggregated statistics from early and late responses to the 45% response-rate Norway in Corona Times survey. Our empirical results suggest that correcting for nonresponse using nonrandom request instruments in subject contact panel data can improve inferences, while posing relatively small challenges for data collection and analysis.
Appendix A Survey Timing
Here we present additional details regarding the implementation of the student entrepreneurship survey that we analyze in Section 3. In each term, survey administrators randomly assigned students into approximately ten strata using a random number generator in Stata. The stratification allowed survey administrators to confirm that request emails were error free for each group prior to sending them to the next group. This feature of the survey administration caused the time-stamps for requests to vary slightly between strata within each term.
[…]
Independence and monotonicity require that responses received in the intervals between time periods are representative and that subjects have sufficient time to respond to each request prior to receiving the next, respectively. Independence could be violated if particular events had systematic effects on self-reported entrepreneurship, which we do not expect for entrepreneurial intention. [^15] One example of a situation in which independence would be threatened by time-varying measurement error is a survey of psychological symptoms with a reminder sent on a holiday. Monotonicity could be violated for some of our requests if some subjects went several days without checking their emails, though these violations would be inconsequential if slow responses indicate relatively high survey aversion among compliers to each request. [^16] Formalizing the relationship between response speed for each request and response aversion exceeds the scope of this paper, but if elapsed time since the first request were a valid instrument used in place of requests, late compliers to request $r$ would have survey aversion between early $r$ compliers and early $r+1$ compliers.
Table A.1: Cumulative Response Rates by Term and Request
Term Request 1 Request 2 Request 3 Request 4 Request 5 Request 6 Request 7
2015, Fall
Request Date 11/23 12/3
Cum. Response Rate 0.16 0.19
2016, Fall
Request Date 10/24 12/12 12/21
Cum. Response Rate 0.11 0.17 0.19
2017, Fall
Request Date 11/7 11/29 12/5
Cum. Response Rate 0.13 0.17 0.19
2018, Fall
Request Date 10/30 11/8 11/29 12/15 12/26
Cum. Response Rate 0.07 0.11 0.12 0.14 0.15
2019, Fall
Request Date 11/1 11/7 12/9 12/21 12/29
Cum. Response Rate 0.07 0.11 0.12 0.12 0.13
2020, Spring
Request Date 4/22 4/27 4/30 5/11
Cum. Response Rate 0.06 0.11 0.14 0.15
2020, Fall
Request Date 10/8 10/15 10/20 11/2 11/17 11/30
Cum. Response Rate 0.08 0.12 0.14 0.16 0.17 0.19
2021, Spring
Request Date 4/21 4/24 4/30 5/11
Cum. Response Rate 0.06 0.10 0.11 0.12
2021, Fall
Request Date 10/20 10/25 11/19 11/30 12/9 12/13 12/28
Cum. Response Rate 0.07 0.11 0.12 0.13 0.14 0.16 0.16
2022, Fall
Request Date 10/27 11/1 11/17 11/29 1/4
Cum. Response Rate 0.04 0.07 0.09 0.09 0.10
Notes: Request timing and cumulative response rates by term.
Appendix B Revisiting the Norway in Corona Times Survey
In this section, we demonstrate the performance of our method on a synthetic version of the Norway in Corona Times (NCT) survey, using only estimated quantities reported by dutz2021selection, henceforth referred to as DHLMTV. [^17] We use the version revised February 2025, which we downloaded from NBER on February 24, 2025. This survey requested information from 10,000 randomly selected adult Norwegians regarding their employment and wages before and after COVID-19 lockdowns and obtained approximately a 50% response rate. DHLMTV merge the survey with administrative data on survey respondents to investigate monthly earnings before lockdown (converted to USD), monthly earnings after lockdown, the share of individuals with an earnings loss above 20%, the share of individuals employed before lockdown, the share of individuals employed after lockdown, and the share of individuals who experienced an employment loss after the lockdown. Importantly for us, these authors provide estimated averages of variables for early and late responders to the NCT survey as well as ground truth values taken from Norwegian administrative data on the same variables.
We begin by matching the moments from DHLMTV’s Table 4. DHLMTV focus on the 93% of subjects who were invited to take the survey online. We further restrict our focus to the 40% of subjects who received no randomized survey incentives to illustrate our method’s performance using only requests as instruments. We therefore construct a synthetic sample of $3,720=10,000\times 0.93\times 0.40$ subjects. Following their description of their methods, we assume that the always taker estimates they report in their Table 4 are consistent estimates for average early responses, and that the reminder complier estimates in their Table 4 are consistent estimates for average late responses.
We produce synthetic data to match DHLMTV Table 4, defining the first 38% of subjects as always-takers, the next 7% as reminder compliers, and the last 55% as nonrespondents. For continuous variables, we first set the means for each group equal to those reported, leaving nonrespondents missing. Next, we then add the standard deviations, backed out from the reported group-specific standard errors, to half of the subjects in each group and subtract them from the other half. For binary variables, we set variable values for always takers and reminder compliers to 1 and 0 as needed to make the group-specific averages match those reported in DHLMTV Table 4. We set variables to missing for the remaining 55% of the observations that do not correspond to always-takers or reminder compliers. We then reshape the data so that each subject has one observation each for $R_{i0}=0$, $R_{i1}=1$, and $R_{i2}=2$, where we initialize $\hat{S}_{it}=0$ for all $i$ and $t$, replaced with $\hat{S}_{i1}=1$ for always-takers and $\hat{S}_{i2}=1$ for both always-takers and reminder compliers.
Table B.2: Methods Comparison using Synthetic Norway in Corona Times Survey
Earnings Employment
Before After Large Loss Before After Loss
Estimator (1) (2) (3) (4) (5) (6)
Panel 1: Ground truth 3,095 2,981 0.148 0.567 0.494 0.091
Panel 2: Sample Means
Always Taker (38%) 3,746 3,783 0.13 0.65 0.64 0.03
(116) (107) (0.01) (0.01) (0.01) (0.00)
Reminder Complier (7%) 3,244 3,257 0.12 0.55 0.55 0.03
(256) (251) (0.02) (0.03) (0.03) (0.01)
Always+Reminders (45%) 3,668 3,701 0.13 0.63 0.63 0.03
(106) (98) (0.00) (0.01) (0.01) (0.00)
Panel 3: DHLMTV (Mid)point 3,368 3,232 0.142 0.588 0.536 0.091 Bounds [0.567, 0.609] Panel 4: Our Method MLE 3,197 3,211 0.116 0.517 0.521 0.026 (250) (238) (0.023) (0.039) (0.039) (0.011) Two step 3,107 3,113 0.116 0.520 0.523 0.026 (303) (297) (0.026) (0.037) (0.037) (0.014) Observations 3,720 3,720 3,720 3,720 3,720 3,720 Notes: Estimates of earnings and employment before and after COVID-19 lockdowns using the NCT survey. Panels 1-3 reconstructed from selected rows of Tables 4 and 5 of Dutz et al (2025). Row 1 of Panel 4 reports full information maximum likelihood estimates of heckman1979sample and van1981demand selection models for continuous and binary variables, respectively, with standard errors clustered at the individual level in parenthesis. Row 2 of Panel 4 reports estimates from the two-step estimator of heckman1979sample, with bootstrapped standard errors with individual-clustered resampling in parentheses.
Our Table B compares estimates using our method with those of Tables 4 and 5 of DHLMTV. Our results in Section 2 regarding the validity of request instruments for identifying local average responses, but do not stipulate a particular means of extrapolation to population averages. For simplicity, we present point estimates obtained from the full information maximum likelihood version of the heckman1979sample selection correction for continuous variables as well as the probit version described by van1981demand for binary variables. We also present estimates from the two-step correction of heckman1979sample for both continuous and binary variables.
We compare estimates of survey variable averages using request instruments to ground truth averages, the pooled means of always-takers and reminder compliers, and DHLMTV’s estimates from their specification reported in their Table 5. DHLMTV’s specification makes use of multiple requests, randomized incentives, and covariate adjustment in the context of a two-dimensional response aversion model. We refer readers to DHLMTV for further details on their assumptions and estimation procedures. To reiterate, our method estimates a one-dimensional response aversion model using multiple requests without using randomized incentives or covariate adjustment. Our method is capable of making use of covariates, but we do not have access to such data for the NCT survey, so we settle for only making use of the estimated means for always takers and reminder compliers for each outcome variable.
Our assessment is that our method performs better than the uncorrected means denoted in Table B as “Always+Reminder.” Specifically, our point estimates are further from the ground truth than the uncorrected estimates for only the Large Loss variable, they are equally distant for the Loss variable, and they are closer for all others. Meanwhile, we obtain larger standard errors which allow us to better avoid the illusion of certainty in cases where point estimates are imperfect. For instance, although our MLE population average estimate for Large Loss is further from the ground truth than the sample mean, we strongly reject that the sample mean is an estimate of the ground truth average, whereas we fail to reject this hypothesis for our preferred estimator. This pattern is repeated for other outcomes, where our preferred method incorrectly rejects the ground truth with 95% confidence only for employment loss, while sample means differ from ground truth averages of all variables with 95% confidence.
Relative to the method of DHLMTV, our method performs somewhat better with continuous variables and substantially worse with binary variables. Specifically, our point estimates are closer to the ground truth than those of DHLMTV for earnings before, earnings after, and employment after, where as their method’s estimates are closer to the ground truth than ours for large earnings losses, employment before, and employment loss. Our method performs particularly poorly with employment loss, which exhibits no variation between always takers and reminder compliers, leading our method to reproduce the in-sample mean as its best guess at the population mean. Meanwhile, DHLMTV’s method obtains the ground truth for this variable exactly.
Our method performs at its worst for estimating a large loss in earnings and employment loss. One explanation for this is that these variables have means near zero, and our normality assumptions may be particularly at odds with reality for such variables. Another explanation is that individuals who have experienced negative outcomes may be particularly resistant to responding to (poorly incentivized) surveys about those outcomes. We are unable to distinguish these explanations with the data at hand, so we end this section by recommending caution to researchers using our method for rare outcomes and/or those that represent negative life outcomes. Our conclusion from this section is that methods that leverage both multiple requests and randomized incentives, such as that of DHLMTV, may perform better than only using multiple requests in some cases, but that our method is preferable to uncorrected estimates in cases where randomized survey response incentives were not administered.

### v4

@@
 We show that if subjects’ accumulated requests and responses over time are recorded and organized as panel data, requests can be used as instruments to correct for nonresponse bias even if total requests are not randomized between subjects.
-We demonstrate our method by estimating an 18-percentage-point gender gap in entrepreneurial career intentions using a survey of undergraduates at the University of Wisconsin-Madison with a 15% response rate and a 20-percentage-point intention gap among respondents.
+We demonstrate our method by estimating an 18-percentage-point gender gap in entrepreneurial career intentions using a survey of undergraduates at the University of Wisconsin-Madison.
 JEL Codes: C83, L26, J16
@@
 This model is identified with two requests, so we also implement and pass a “predictable trends” overidentification test of the null that later requests have no direct effect on the outcome, supporting the validity of our parametric and nonparametric assumptions in this application.
-In Appendix B, we estimate labor market outcomes using a synthetic version of the Norway in Corona Times survey.
-We construct our synthetic dataset to match respondent mean estimates reported by dutz2021selection, which are sufficient to estimate population means with our method.
+In Supplemental Appendix B, we estimate labor market outcomes using a synthetic version of the Norway in Corona Times survey constructed to match estimates reported by dutz2021selection.
 The 95% confidence intervals of our estimated population means cover ground truth averages from administrative data reported by these authors for five out of six survey variables, with respondent averages differing significantly from the ground truth for all variables.
@@
 [^7] We use this definition in lieu of $S_{it}=S_{it}(R_{it})$ because we assume that subjects respond at most once.
-Similarly, let $Y_{it}=Y_{it}(S_{it},R_{it})=Y_{it}(1,R_{it})S_{it}+Y_{it}(0,R_{it})(1-S_{it})$ with $Y_{it}(s,r)$ denoting the response subject $i$ would give at time $t$ if they made response choice $s$ given $r$ accumulated requests, with $Y_{it}(0,r)$ missing for all $i$, $t$, and $r$.
-$S_{it}(r)$ and $Y_{it}(s,r)$ are defined for all values in the supports of their domains, $r\in\{0,...,\max(R_{it})\}$ and $s\in\{0,1\}$, respectively, while they are observed only for their realized values.
+Similarly, let $Y_{it}=Y_{it}(S_{it},R_{it})=Y_{it}(1,R_{it})S_{it}+Y_{it}(0,R_{it})(1-S_{it})$ with $Y_{it}(1,r)$ denoting the response subject $i$ would give at time $t$ if were to respond given $r$ accumulated requests, with $Y_{it}(0,r)$ missing for all $i$, $t$, and $r$.
+$S_{it}(r)$ and $Y_{it}(1,r)$ are defined for all $r\in\{0,...,\max(R_{it})\}$, respectively, while they are observed only for the realized $R_{it}$.
 Following the marginal treatment response framework of heckman2005structural; heckman2007econometric, we assume that selection bias is driven by dependence between responses and response aversion.
@@
 It follows that we can learn about $m(u)$ if we can estimate $\mathbb{E}[Y_{i}^{*}|S_{it}(r)-S_{it}(r^{\prime})=1]$, which we refer to as the local average response for $r-r^{\prime}$ compliers.
-If $m(u)$ is parameterized, as in a heckman1979sample model, point estimation is possible if there are weakly more identified values of $P(r)$ than parameters.
+If $m(u)$ is parameterized, as in a heckman1979sample model, point estimation is possible if there are weakly more identified local average responses than parameters.
 Otherwise, $m(u)$ can be bounded using methods such as those of horowitz2000nonparametric or lee2009training.
@@
 […]
-[…]
+This assumption strengthens the independence assumption of imbens1994identification by disallowing requests from being selectively sent to subjects with nonrepresentative potential responses to that request or any prior request.Randomizing requests and setting $T=1$, as proposed by dinardo2021practical, supports this assumption, with it also holding for $T>1$ with uniform requests if potential responses and response choices are time-invariant on average.
+Less extreme coarsening of the time index, for instance defining $t$ in weeks instead of minutes, can support this assumption by essentially averaging over temporal idiosyncrasies that occur in continuous time.
 We next assume that requests do not directly affect potential responses, allowing us to write $Y_{it}(1)=Y_{it}(1,r)$.
@@
 In contrast with intertemporal independence, exclusion is more likely to be violated due to continuous time variation in potential responses and requests if time periods are defined coarsely, as subjects receiving different numbers of requests in the same discrete time period may receive them at different points in continuous time.
-[^9] Following imbens1994identification, we’ve assumed implicitly that $Y_{it}(s,r)$ and $S_{it}(r)$ are fixed for all $s,r$, for all $i,t$ (SUTVA).
-This assumption is consequentially violated if potential responses or potential response choices systematically vary in continuous time within discrete time periods.
 With independence and exclusion, differences in responses between requests can only be explained by differences in responses between distinct complier groups, defined by common values of $S_{it}(r)$ for all $r$.
@@
 $\mathbb{E}[Y_{it}(1,r)S_{it}(r)]=\mathbb{E}[Y_{i}^{*}S_{it}(r)]$ for all $r$.
-This assumption is weaker than $Y_{it}(1,r)=Y_{i}^{*}$, which is common in the related literature, because it allows for idiosyncratic shocks that cancel out within complier groups, such as if students’ potential responses regarding career intentions change over time in response to new information, such as grades on assignments.
-[^10] We hope that our consideration of measurement error will forestall inaccurate presumptions that weaker assumptions, such as $\mathbb{E}[Y_{it}(1,r)]=\mathbb{E}[Y_{i}^{*}]$, are sufficient.
-Intertemporal independence and exclusion rule out systematic time-varying measurement error and measurement error caused by requests, respectively, but they do not rule out systematic inaccuracies within or across complier groups.
+This assumption is weaker than $Y_{it}(1,r)=Y_{i}^{*}$, which is common in the related literature, because it allows for idiosyncratic shocks that cancel out within complier groups, such as if students’ potential responses regarding career intentions change over time in response to new information, such as grades on assignments.Intertemporal independence and exclusion rule out systematic time-varying measurement error and measurement error caused by requests, respectively, but they do not rule out systematic inaccuracies within or across complier groups.
 For example, if less conscientious subjects give systematically inaccurate responses to late requests, nonresponse corrections using request instruments will tend to replicate similar inaccuracy in their predictions for non-respondents.
@@
 This application showcases the importance of researcher judgments in defining the time index, adjusting for nonrandom variation in requests, and testing modeling assumptions in cases where there are more requests than are needed for identification.
-In Appendix B, we demonstrate our method’s accuracy by showing that nonresponse-corrected estimates are closer to ground truth values from administrative data than respondent averages for the Norway in Corona Times Survey, using survey and ground truth averages reported by dutz2021selection.
+In Supplemental Appendix B, we demonstrate our method’s accuracy by showing that nonresponse-corrected estimates are closer to ground truth values from administrative data than respondent averages for the Norway in Corona Times Survey, using survey and ground truth averages reported by dutz2021selection.
 3.1 Data
@@
 We code an “Intention” variable equal to 1 if the student answered “Yes” and 0 if they answered “No” or “I Don’t Know.” We observe term-specific survey request stratum indicators for groups of students who are sent requests at the same time through survey software.
-We observe timestamps, in seconds, for each request made to each stratum and for each student’s response if they respond.
+We observe timestamps for each request made to each stratum and for each student’s response if they respond.
 Students who opted out of the survey in a prior year have no stratum indicators and no request or response timestamps in the raw data, whereas students who respond to early requests have stratum indicators and request timestamps for all intended requests in raw data, though they did not receive additional actual requests after responding.
@@
 3.2 Constructing a Request-Response Panel
-Our results from Section 2 require panel data wherein subjects are observed multiple times, so we construct such a panel.
+Our results from Section 2 involve panel data wherein subjects are observed multiple times as data requests are made, so we construct such a panel for each term.
+We index terms/semesters by $s$, with retained responses, retained response choices, and accumulated requests denoted by $\hat{Y}_{ist}$, $\hat{S}_{ist}$, and $R_{ist}$, respectively.
 In our setting, students who opted out of the survey in a prior term and students who respond early in a given term do not receive subsequent requests.
 It follows that actual requests almost surely violate the independence assumption in Section 2, as students receiving each number of requests differ systematically.
-We address this by randomly assigning opt-outs to strata ex post and assuming $S_{i}(r)=0$ for all $r$ for these students, then imputing intended request timestamps for each student using the request timestamps of nonrespondents in their stratum.
-For each student-term in a stratum, we produce observations for $t$ in $0,1,...,R_{i}$ such that time periods are defined as intervals in which a subject has received a given number of intended requests, with $R_{it}=t$ for all $i$ and $t$.
+We address this by randomly assigning opt-outs to strata ex post and assuming $S_{ist}(r)=0$ for all $t$ and all $r$ for such students, then imputing intended request timestamps for each student using the request timestamps of nonrespondents in their stratum.
+For each student-term in a stratum, we produce observations for $t$ in $0,1,...,R_{i}$ such that time periods are defined as intervals in which a subject has received a given number of intended requests, with $R_{ist}=t$ for all $i$, $s$, and $t$.
 Independence and monotonicity require that the times at which requests were made had representative potential responses and that subjects had sufficient time to respond between requests, respectively.
 We report request timestamps in Appendix A, revealing that the shortest interval between requests is three days, which we assume was adequate time for students to respond.
-We set $S_{it}=1$ if subject $i$’s response timestamp is between those defining the beginning and end of period $t$, with $Y_{it}$ given by their response to the survey if $S_{it}=1$.
-We forward-impute retained response choices $\hat{S}_{it}$ and retained responses $\hat{Y}_{it}$ as described in Section 2.
+We set $S_{ist}=1$ if subject $i$’s response timestamp is between those defining the beginning and end of period $t$, with $Y_{ist}$ given by their response to the survey if $S_{ist}=1$.
+We forward-impute retained response choices $\hat{S}_{ist}$ and retained responses $\hat{Y}_{ist}$ as described in Section 2 for each value of $s$.
 3.3 Nonresponse-Corrected Entrepreneurial Intentions by Gender
@@
 Because requests vary in number from two to seven between terms, we drop all observations with $R_{it}>2$ to support our independence assumption for the current specification.
-We estimate a parametric van1981demand selection model separately by gender via maximum likelihood, in which
-$\hat{S}_{it}=\begin{cases}&1\mbox{ if }Z_{it}\alpha+u_{i}\geq 0,\\ &0\mbox{ otherwise,}\end{cases}$ (4)
-$\hat{Y}_{it}=\begin{cases}1\mbox{ if }X_{i}\beta+\epsilon_{i}\geq 0\quad\&\quad&\hat{S}_{it}=1,\\ 0\mbox{ if }X_{i}\beta+\epsilon_{i}<0\quad\&\quad&\hat{S}_{it}=1,\\ \cdot\mbox{ if }&\hat{S}_{it}=0,\end{cases}$ (5)
+We estimate a parametric van1981demand selection model via maximum likelihood, separately for men and women.
+Specifically, we assume
+$\hat{S}_{ist}=\begin{cases}&1\mbox{ if }Z_{ist}\alpha+u_{ist}\geq 0,\\ &0\mbox{ otherwise,}\end{cases}$ (4)
+$\hat{Y}_{ist}=\begin{cases}1\mbox{ if }X_{is}\beta+\epsilon_{ist}\geq 0\quad\&\quad&\hat{S}_{ist}=1,\\ 0\mbox{ if }X_{is}\beta+\epsilon_{ist}<0\quad\&\quad&\hat{S}_{ist}=1,\\ 0\mbox{ if }&\hat{S}_{ist}=0,\end{cases}$ (5)
 and
 […]
-where $X_{i}=1$, $Z_{it}=[X_{i},I(R_{it}=2)]$, and $\rho$ gives the correlation between outcome residuals $\epsilon_{i}$ and response preference $u_{i}=\Phi^{-1}(1-U_{i})$, where $U_{i}$ is $i$’s response aversion percentile in the notation of Section 2 and $\Phi^{-1}(\cdot)$ denotes the inverse of the normal CDF.
+where $X_{is}=1$, and $Z_{ist}=[X_{is},I(R_{ist}=2)]$.
+We assume $(\epsilon_{ist},u_{ist})$ are independent and identically distributed between $i$, with potential dependence allowed between $s$ and $t$ for each $i$.
+$\rho$ gives the correlation between outcome residuals $\epsilon_{ist}$ and response preference $u_{ist}=\Phi^{-1}(1-U_{ist})$, where $U_{ist}$ is $i$’s response aversion percentile in the notation of Section 2 and $\Phi^{-1}(\cdot)$ denotes the inverse of the normal CDF.
 [^9] Our 15% response rate is insufficient for informative bounds from conservative approaches such as that of horowitz2000nonparametric, and we doubt that gender satisfies the monotonicity assumption of lee2009training or behaghel2015please, rendering the “local gender gap” among survey respondents unidentified, to say nothing of its policy-relevance.
-This model implies that expected intention as a function of response aversion is $m(u)=\Phi((X_{i}\beta+\rho\Phi^{-1}(1-u))/\sqrt{1-\rho^{2}})$, with the expectation conditional only on $X_{i}$ given by $\int_{0}^{1}m(u)du=\Phi(X_{i}\beta)$.
+This gender-specific constant-only model implies that expected intention as a function of response aversion is $m(u)=\Phi((\beta+\rho\Phi^{-1}(1-u))/\sqrt{1-\rho^{2}})$, with the unconditional expectation expressed as $\int_{0}^{1}m(u)du=\Phi(\beta)$.
 We graphically present our estimates of $m(u)$ for men and women in Figure 1 for $u\in[0,1]$.
@@
 To implement the decomposition, we estimate the model in (4)-(6) jointly for men and women via maximum likelihood using gender-specific parameters $(\beta_{m},\alpha_{m},\rho_{m})$ for men and $(\beta_{w},\alpha_{w},\rho_{w})$ for women.
-To avoid overweighting terms in which more requests were sent, which have more observations, we weight each observation by the inverse of the total number of requests sent to that student in that term.
-We define $X_{i}$ as all covariates listed in Table 1 along with term fixed effects, and we define $Z_{it}$ as binary indicators for each value in $R_{it}=1,2...,7$ interacted with $X_{i}$.131313blandhol2022tsls recommend rich covariate interactions for instrumental variables when estimating local average treatment effects via two stage least squares.
+To avoid overweighting terms in which more requests were sent, we weight each observation by the inverse of the total number of requests sent to that student in that term.
+We define $X_{is}$ as all covariates listed in Table 1 along with term fixed effects, and we define $Z_{ist}$ as binary indicators for each value in $R_{ist}=1,2...,7$ interacted with $X_{is}$.101010blandhol2022tsls recommend rich covariate interactions for instrumental variables when estimating local average treatment effects via two stage least squares.
 The estimator and target parameters are different here, but we find their argument convincing that if model assumptions hold conditional on $X_{i}$, then $X_{i}$ should be flexibly interacted with instruments to support identification.
@@
 Columns (3) and (6) give uncorrected and corrected estimated contributions of each variable or coefficient to the estimated gap in entrepreneurial intentions, using equations (7), (8), (9), (10), and (11).
-$\mathbb{E}[\mathbb{E}[Y_{i}^{*}|S_{i}=1,X_{i}]]$ is estimated as the average of $\Phi(X_{i}\hat{\beta}^{uncorrected})$ and $\mathbb{E}[\mathbb{E}[Y_{i}^{*}|X_{i}]]$ is estimated as the average of $\Phi(X_{i}\hat{\beta}^{corrected})$ over student-term observations.
+$\mathbb{E}[\mathbb{E}[Y_{is}^{*}|S_{isT}=1,X_{i}]]$ is estimated as the average of $\Phi(X_{is}\hat{\beta}^{uncorrected})$ and $\mathbb{E}[\mathbb{E}[Y_{is}^{*}|X_{is}]]$ is estimated as the average of $\Phi(X_{is}\hat{\beta}^{corrected})$ over student-term observations.
 Term fixed effects, included in $X_{i}$ to support the independence assumption, are treated as part of the unexplained gap.
 Standard errors clustered at the individual level are in parentheses.
-We estimate the contribution of $x_{i}\subset X_{i}$ to the entrepreneurial intention gap as
-$\hat{\Delta}_{x}=\sum_{i=1}^{N_{w}}\frac{\Phi(X_{i}\hat{\beta}_{w}+(\bar{x}_{m}-\bar{x}_{w})\hat{\beta}_{w,x})-\Phi(X_{i}\hat{\beta}_{w})}{N_{w}},$ (7)
+We estimate the contribution of $x_{is}\subset X_{is}$ to the entrepreneurial intention gap as
+$\hat{\Delta}_{x}=\sum_{i=1}^{N_{w}}\frac{\Phi(X_{is}\hat{\beta}_{w}+(\bar{x}_{m}-\bar{x}_{w})\hat{\beta}_{w,x})-\Phi(X_{is}\hat{\beta}_{w})}{N_{w}},$ (7)
 wherein we denote the sample average of $x_{i}$ for men and women as $\bar{x}_{m}$ and $\bar{x}_{w}$, respectively, and we denote male and female coefficients on $x_{i}$ as $\beta_{m,x}$ and $\beta_{w,x}$, respectively.
-[^14] Fairlie’s (fairlie2005extension) decomposition for nonlinear models considers the case in which the marginal distribution of $x_{i}$ for women is set equal to men’s, while ours considers the case in which the mean of $x_{i}$ for women is set equal to men’s via a level shift of women’s distribution.
-We view the counterfactual we consider as likely simpler to understand and simpler to implement for policymakers, though describing $\beta_{w,x}$ as the “effect” of changing $x_{i}$ for women requires assumptions beyond those that either we or fairlie2005extension make.
-We similarly estimate the contribution to the average gender gap of $x_{i}$’s differential effect on $Y_{i}$ between men and women as
-$\hat{\Delta}_{\beta_{x}}=\sum_{i=1}^{N_{w}}\frac{\Phi(X_{i}\hat{\beta}_{w}+x_{i}(\hat{\beta}_{m,x}-\hat{\beta}_{w,x}))-\Phi(X_{i}\hat{\beta}_{w})}{N_{w}}.$ (8)
-In addition to estimating the contribution of each variable and its coefficient on the gender-intention gap, we also estimate the contribution of the entire vector $X_{i}$ on the gap as
-$\hat{\Delta}_{X}=\sum_{i=1}^{N_{w}}\frac{\Phi(X_{i}\hat{\beta}_{w}+(\bar{X}_{m}-\bar{x}_{w})\hat{\beta}_{w})-\Phi(X_{i}\hat{\beta}_{w})}{N_{w}},$ (9)
-defining the vector of the means of all elements in $X$ for people with $W_{i}=w$ as $\bar{X}_{w}$.
+[^11] Fairlie’s (fairlie2005extension) decomposition for nonlinear models considers the case in which the marginal distribution of $x_{is}$ for women is set equal to men’s, while ours considers the case in which the mean of $x_{is}$ for women is set equal to men’s via a level shift of women’s distribution.
+We view the counterfactual we consider as likely simpler to understand and simpler to implement for policymakers, though describing $\beta_{w,x}$ as the “effect” of changing $x_{is}$ for women requires assumptions beyond those that either we or fairlie2005extension make.
+We similarly estimate the contribution to the average gender gap of $x_{is}$’s differential effect on $Y^{*}_{is}$ between men and women as
+$\hat{\Delta}_{\beta_{x}}=\sum_{i=1}^{N_{w}}\frac{\Phi(X_{is}\hat{\beta}_{w}+x_{is}(\hat{\beta}_{m,x}-\hat{\beta}_{w,x}))-\Phi(X_{is}\hat{\beta}_{w})}{N_{w}}.$ (8)
+In addition to estimating the contribution of each variable and its coefficient on the gender-intention gap, we also estimate the contribution of the entire vector $X_{is}$ on the gap as
+$\hat{\Delta}_{X}=\sum_{i=1}^{N_{w}}\frac{\Phi(X_{is}\hat{\beta}_{w}+(\bar{X}_{m}-\bar{X}_{w})\hat{\beta}_{w})-\Phi(X_{is}\hat{\beta}_{w})}{N_{w}},$ (9)
+defining the vector of sample means of all covariates in $X_{is}$ for men as $\bar{X}_{m}$ and for all women as $\bar{X}_{w}$.
 Similarly, we estimate the contribution of the entire vector of gender differences in coefficients on the gender-intention gap as
-$\hat{\Delta}_{\beta}=\sum_{i=1}^{N_{w}}\frac{\Phi(X_{i}\hat{\beta}_{w}+X_{i}(\hat{\beta}_{m}-\hat{\beta}_{w}))-\Phi(X_{i}\hat{\beta}_{w})}{N_{w}}.$ (10)
+$\hat{\Delta}_{\beta}=\sum_{i=1}^{N_{w}}\frac{\Phi(X_{is}\hat{\beta}_{w}+X_{is}(\hat{\beta}_{m}-\hat{\beta}_{w}))-\Phi(X_{is}\hat{\beta}_{w})}{N_{w}}.$ (10)
 We estimate the remaining gap, which is driven by the interaction of male-female differences in $X$ and male-female differences in $\beta$, distributional differences in $X$ between men and women, and nonlinearity of the normal distribution as
-$\hat{\Delta}_{R}=\sum_{i=N_{w}+1}^{N}\frac{\Phi(X_{i}\hat{\beta}_{m})}{N_{m}}-\sum_{i=1}^{N_{w}}\frac{\Phi(X_{i}\hat{\beta}_{w})}{N_{w}}-\hat{\Delta}_{X}-\hat{\Delta}_{\beta}.$ (11)
+$\hat{\Delta}_{R}=\sum_{i=N_{w}+1}^{N}\frac{\Phi(X_{is}\hat{\beta}_{m})}{N_{m}}-\sum_{i=1}^{N_{w}}\frac{\Phi(X_{is}\hat{\beta}_{w})}{N_{w}}-\hat{\Delta}_{X}-\hat{\Delta}_{\beta}.$ (11)
 Table 2 contains estimated means of the variable or coefficient listed in each row for men, women, and the contribution of that item to the intention gap, using predicted values from a naive probit in columns 1–3, and using our proposed selection correction model in columns 4–6.
-$X$ is observed for all individuals regardless of survey response, so the means of $X$ by gender are the same for the uncorrected model and the corrected model.
+$X_{is}$ is observed for all individuals regardless of survey response, so the means of $X_{is}$ by gender are the same for the uncorrected model and the corrected model.
 Columns 3 and 6 report estimated contributions of each row quantity to the gender gap as described in equations (7), (8), (9), (10), and (11).
-We perform our decomposition for all variables in Table 1, with term fixed effects included in $X_{i}$ contributing only to the unexplained gap.
+We perform our decomposition for all variables in Table 1, with term fixed effects included in $X_{is}$ contributing only to the unexplained gap.
 We estimate the overall gender gap by calculating
-$\hat{\Delta}=\sum_{i=N_{w}+1}^{N}\frac{\Phi(X_{i}\hat{\beta}_{m})}{N_{m}}-\sum_{i=1}^{N_{w}}\frac{\Phi(X_{i}\hat{\beta}_{w})}{N_{w}},$ (12)
+$\hat{\Delta}=\sum_{i=N_{w}+1}^{N}\frac{\Phi(X_{is}\hat{\beta}_{m})}{N_{m}}-\sum_{i=1}^{N_{w}}\frac{\Phi(X_{is}\hat{\beta}_{w})}{N_{w}},$ (12)
 where our estimates of $(\hat{\beta}_{w},\hat{\beta}_{m})$ differ by estimation method.
 The estimated gender-intention gap from our preferred specification is 18.3 percentage points, with a naive probit producing an estimate of 20.2 percentage points.
-Differences between the corrected and uncorrected model are due to the inclusion of the $\rho$ parameter in the corrected model (with $\rho=0$ fixed in the uncorrected model) which is positive if early respondents have higher intention and negative if they have lower intention.
+Differences between the corrected and uncorrected model are due to the use of only observations with $t=T$ for the naive probit and the inclusion of the $\rho$ parameter in the corrected model (with $\rho=0$ fixed in the uncorrected model) which is positive if early respondents have higher intention and negative if they have lower intention.
 Our estimates of $\rho$ in the current specification are insignificant, suggesting limited nonresponse bias, though we see the same pattern with the signs of our point estimates here as we did in the no-controls specification shown in Figure 1.
@@
 The difference in corrected estimates in Section 3.3 and the corrected estimates in Table 2 is due to the inclusion of controls and observations for requests three through seven, which were observed in some years but not others.
-The richer nonresponse correction specification in this section approximates a nonparametric specification that would fit a curve through local responses as in Figure 1 for each value of $X_{i}$, where the slope of the curve is determined by a gender-specific $\rho$ that is held fixed across $X_{i}$.
+The richer nonresponse correction specification in this section approximates a nonparametric specification that would fit a curve through local responses as in Figure 1 for each value of $X_{is}$, where the slope of the curve is determined by a gender-specific $\rho$ that is held fixed across $X_{i}$.
 For both men and women, $\rho$ is attenuated in the richer model, suggesting that observed covariates account for both late responses and differences in intention between late and early responses.
-Nonetheless, the same general pattern from Figure 1 persists, with positive selection for men ($\hat{\rho}_{m}>0)$ and negative selection for women $(\hat{\rho}_{f}<0)$, though these estimates are not statistically significant in this specification.
+Nonetheless, the same general pattern from Figure 1 persists, with positive selection for men ($\hat{\rho}_{m}>0)$ and negative selection for women $(\hat{\rho}_{w}<0)$, though these estimates are not statistically significant in this specification.
 Our estimates are only reliable if the assumptions in Section 2 and equation (6) hold for this application.
@@
 To do so, we construct a panel dataset that contains a row for each student for each data request they receive, with their accumulated requests recorded in all rows and their responses recorded only in rows representing time periods after they respond.
-We find a large gender gap both without correcting for nonresponse and when correcting for nonresponse using requests as instruments in a van1981demand model, with the uncorrected estimates strongly rejecting the point estimate of the average gender gap obtained when correcting for nonresponse.
-In an additional application in Appendix B, we accurately and precisely estimate ground truth averages from administrative data for five out of six labor market indicators using only aggregated statistics from early and late responses to the 45% response-rate Norway in Corona Times survey.
+We find a large gender gap uncorrected estimates strongly rejecting the point estimate of the average gender gap obtained when correcting for nonresponse.
+In an additional application in Supplemental Appendix B, we accurately and precisely estimate ground truth averages from administrative data for five out of six labor market indicators using only aggregated statistics from early and late responses to the 45% response-rate Norway in Corona Times survey.
 Our empirical results suggest that correcting for nonresponse using nonrandom request instruments in subject contact panel data can improve inferences, while posing relatively small challenges for data collection and analysis.